# OBLR-PO: A Theoretical Framework for Stable Reinforcement Learning

### v1

@@ 1 Introduction
Reinforcement learning (RL) has become a central paradigm for training large-scale models to exhibit complex reasoning and decision-making abilities [christiano2023deepreinforcementlearninghuman, stiennon2022learningsummarizehumanfeedback, ouyang2022traininglanguagemodelsfollow, bai2022traininghelpfulharmlessassistant, deepseekai2025deepseekr1incentivizingreasoningcapability, sheng2025solving]. Policy optimization algorithms such as Proximal Policy Optimization (PPO) [schulman2017proximalpolicyoptimizationalgorithms], Group Relative Policy Optimization (GRPO) [deepseekai2025deepseekr1incentivizingreasoningcapability, shao2024deepseekmathpushinglimitsmathematical], and REINFORCE with leave-one-out baselines (RLOO) [ahmadian2024basicsrevisitingreinforcestyle] explore different designs of reward modeling and advantage estimation to guide learning. Despite these advances, training stability remains a fundamental bottleneck, particularly in the post-training stage of large language models [castanyer2025stablegradientsstablelearning, roux2025taperedoffpolicyreinforcestable, zhang2025r1rewardtrainingmultimodalreward, su2025trustregionpreferenceapproximation].
Training stability is influenced not only by algorithmic design but also by choices such as learning rate schedules and baseline functions. In large-scale pre-training, adaptive or decayed learning rate schedules are routinely used to ensure stable optimization and steady convergence [bengio2012practicalrecommendationsgradientbasedtraining, bergsma2025straightzerolinearlydecaying, darken1990note, li2019exponential, darken1992learning, li2025unveilingrolelearningrate]. It is therefore natural to ask whether similar benefits could be realized in reinforcement learning, where instability remains a persistent challenge [wang2025octothinkermidtrainingincentivizesreinforcement]. Baseline design has likewise been explored as a practical tool for variance reduction, with numerous heuristics proposed across different policy optimization methods [zhu2023finetuninglanguagemodelsadvantageinduced, ahmadian2024basicsrevisitingreinforcestyle, walder2025passkpolicyoptimizationsolving]. However, most of these strategies are empirical in nature, and the field still lacks systematic analysis that clarifies what constitutes an effective baseline and how it influences learning dynamics.
These limitations highlight the need for deeper theoretical guidance. While empirical techniques for stability are widely adopted, systematic theory for post-training policy optimization is still scarce [pang2025theorypracticegrpotrajectorycorrected, zhang2025speedrlfastertrainingreasoning, brantley2025acceleratingrlllmreasoning, yao2025optimizingchainofthoughtreasonersgradient, li2025stayingsweetspotresponsive]. Under mild and interpretable assumptions, rigorous analysis can illuminate the statistical properties of gradient estimators, reveal how baseline choices interact with learning dynamics, and clarify when adaptive learning rates can provably improve stability. Such theoretical insights not only fill a long-standing gap in understanding but also directly motivate algorithmic designs that bridge principled analysis with practical effectiveness.
@@ 2 Related Work
Recent studies have analyzed the training dynamics of policy optimization, especially focusing on the loss function upper bound and the resulting convergence guarantees [li2025stayingsweetspotresponsive, pang2025theorypracticegrpotrajectorycorrected, brantley2025acceleratingrlllmreasoning, yao2025optimizingchainofthoughtreasonersgradient]. For example, the impact of a single-step update on the loss, along with the identification of the optimal update vector, has been analyzed [li2025stayingsweetspotresponsive]. Similarly, stochastic no-regret oracle frameworks have been employed to provide theoretical guarantees, leading to regret upper bounds and formal connections to online learning [brantley2025acceleratingrlllmreasoning]. In parallel, classical gradient descent optimization theory has been applied to establish convergence rates under smoothness and convexity assumptions [pang2025theorypracticegrpotrajectorycorrected, yao2025optimizingchainofthoughtreasonersgradient]. Related efforts also bridge theory and practice by providing convergence results for GRPO and related algorithms, thereby supporting their empirical success [pang2025theorypracticegrpotrajectorycorrected]. Moreover, under smoothness conditions, it has been shown that the loss function admits a guaranteed decreasing rate [yao2025optimizingchainofthoughtreasonersgradient]. Building on this analysis, our work establishes tighter upper bounds, provides stronger and more general convergence guarantees, and introduces principled strategies for learning rate schedules and baseline design.
@@
Policy optimization, originating from reinforcement learning, is central to shaping the reasoning capabilities of large language models and has therefore become the foundation of reinforcement learning from human feedback (RLHF) [christiano2023deepreinforcementlearninghuman, stiennon2022learningsummarizehumanfeedback, ouyang2022traininglanguagemodelsfollow, bai2022traininghelpfulharmlessassistant, deepseekai2025deepseekr1incentivizingreasoningcapability]. Within this framework, supervised fine-tuning (SFT) [ouyang2022traininglanguagemodelsfollow, wang-etal-2023-self-instruct] provides initial alignment through imitation on instruction data, PPO [schulman2017proximalpolicyoptimizationalgorithms] extends this paradigm with critic-based reinforcement learning on preference-model rewards, and Direct Preference Optimization (DPO) [rafailov2024directpreferenceoptimizationlanguage] further simplifies the objective by introducing a contrastive loss that directly matches human preferences between responses. To mitigate the complexity of critic-based methods, GRPO [deepseekai2025deepseekr1incentivizingreasoningcapability, shao2024deepseekmathpushinglimitsmathematical] replaces the learned value function with a group-based baseline defined as the average reward within each candidate set, while alternative designs such as ReMax [li2024remaxsimpleeffectiveefficient], RLOO [ahmadian2024basicsrevisitingreinforcestyle], and Reinforce++ [hu2025reinforceefficientrlhfalgorithm] adopt maximum-reward, leave-one-out, or variance-reduced baselines, respectively. Building on these developments, this work presents a general formulation that unifies the above algorithms under a common framework, enabling systematic theoretical analysis and leading to the identification of a principled optimal baseline.
@@ 3.2 Related RL Algorithms
The Proximal Policy Optimization (PPO) [schulman2017proximalpolicyoptimizationalgorithms] objective is formulated as
@@
where the advantage estimator $\hat{A}^{\operatorname{PPO}}$ is computed using Generalized Advantage Estimation (GAE) [schulman2018highdimensionalcontinuouscontrolusing], a widely adopted variance-reduction technique that stabilizes policy-gradient training.
@@
Group Relative Policy Optimization (GRPO) [shao2024deepseekmathpushinglimitsmathematical] optimizes policies by comparing rewards among a group of sampled outputs, without relying on an explicit value or reward model. Given a query $q$, we draw $G$ outputs $\{o_{i}\}_{i=1}^{G}$ from the old policy $\pi_{\theta_{\operatorname{old}}}$ with associated rewards $r_{i}=F(q,o_{i})$. The objective is
@@
The ReMax [li2024remaxsimpleeffectiveefficient] method draws inspiration from the REINFORCE with Baseline approach, where we modify the gradient estimation by incorporating a subtractive baseline value. The objective is:
@@
REINFORCE Leave-One-Out (RLOO) [ahmadian2024basicsrevisitingreinforcestyle, Kool2019Buy4R] extends the REINFORCE estimator to the multi-sample setting by employing a leave-one-out baseline. Given a query $q$, we draw $G$ outputs $\{o_{i}\}_{i=1}^{G}$ from the old policy $\pi_{\theta_{\operatorname{old}}}$ with rewards $r_{i}=F(q,o_{i})$. The objective is
@@ 3.4 Assumptions
PPO [schulman2017proximalpolicyoptimizationalgorithms]
GRPO [shao2024deepseekmathpushinglimitsmathematical]
ReMax [li2024remaxsimpleeffectiveefficient]
RLOO [ahmadian2024basicsrevisitingreinforcestyle]
@@ 4.1 Bias and Variance Analysis
Next, we analyze the covariance of the noise term $\xi(\theta)$. To simplify the expression, we define $\mathbf{H}(\theta)$ as the covariance matrix with respect to the policy $\pi_{\theta}$, i.e.,
$\boldsymbol{H}(\theta):=\text{Cov}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)(F(q,o)-b_{\theta}(q))\right]$ (25)
Theorem 2 (Covariance Expression).
@@
$\text{Cov}[\xi(\theta)]=\frac{1}{N_{t}G_{t}}\boldsymbol{H}(\theta)$ (26)
@@ 4.2 Deriving an Upper Bound for the Loss Function
$\mathbb{E}[\mathscr{L}(\theta_{T})]$ $\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}$
$\quad+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}$
$\quad+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\left(\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$ (28)
@@ 4.3 Optimal Learning Rate Schedule
$\eta_{t}$ $=\frac{1}{BL+B^{2}M}\cdot\frac{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}$ (30)
$=\frac{1}{BL+B^{2}M}\cdot\frac{N_{t}G_{t}\operatorname{SNR}(\theta_{t})}{1+N_{t}G_{t}\operatorname{SNR}(\theta_{t})}.$ (31)
@@
$\operatorname{SNR}(\theta)$ $=\frac{\|\nabla_{\theta}[J(\theta)]\|_{2}^{2}}{\mathbb{E}[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\,\hat{A}^{\mathrm{PO}}(q,o)-\nabla_{\theta}[J(\theta)]\|_{2}^{2}]}$ (32)
$=\frac{\|\nabla_{\theta}\mathscr{L}(\theta)\|_{2}^{2}}{\operatorname{tr}(\boldsymbol{H}(\theta))}.$ (33)
This theorem shows that the optimal learning rate is governed by the signal-to-noise ratio (SNR) of the gradient [zhang2025speedrlfastertrainingreasoning, he2025deltalnormalizationrethink] and is further amplified as data broadens ($N_{t}$) and knowledge deepens ($G_{t}$[yang2025depthbreadthsynergyrlvrunlocking]. The following two takeaways highlight these messages.
@@
Takeaway 2 As $N_{t}$ broadens and $G_{t}$ deepens knowledge, the learning rate should rise accordingly.
@@
$-\sum_{t=0}^{T-1}\frac{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{2(BL+B^{2}M)\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\text{tr}(\mathbf{H}(\theta_{t}))\right)}.$ (34)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}=O\left(\frac{1}{\sqrt{T}}\right),$ (35)
@@ 4.4 Optimal Baseline Design
This expression reveals that the optimal baseline depends on the gradient magnitudes rather than uniform averaging, leading to the following insight.
Takeaway 3 The baseline should not be a simple average, but a gradient-weighted one.
@@ 5 Methodology
$\hat{\eta}_{t}=\eta_{0}\frac{N_{t}G_{t}\widehat{\operatorname{SNR}(\theta_{t})}}{1+N_{t}G_{t}\widehat{\operatorname{SNR}(\theta_{t})}}$, (37)
@@ 6 Experiments
In this section, we present a comprehensive overview of our experimental setup, evaluation metrics, and results. We evaluate the proposed Optimal Baseline and Learning-Rate Policy Optimization (OBLR-PO) algorithm on the Qwen3-4B-Base and Qwen3-8B-Base models [yang2025qwen3technicalreport], comparing its performance against the Group Relative Policy Optimization (GRPO) baseline. All experiments are implemented with the VERL framework [sheng2024hybridflow] and conducted on four H200 GPUs to enable large-scale training. The subsequent subsections detail the training configuration, evaluation protocols, and empirical results.
@@
  • OlympiadBench [he2024olympiadbench]: A benchmark of olympiad-level bilingual multimodal scientific problems, designed to evaluate high-difficulty reasoning, cross-modal understanding, and advanced problem-solving capabilities of language models.
  • GSM8K [cobbe2021trainingverifierssolvemath]: A benchmark of grade-school level math word problems, designed to measure arithmetic reasoning and step-by-step problem-solving skills of language models.
  • AIME25 [aime25]: A benchmark consisting of problems from the 2025 American Invitational Mathematics Examination (AIME), designed to assess advanced mathematical reasoning, precise numeric computation, and multi-step problem-solving abilities of language models.
  • MATH500 [hendrycksmath2021, lightman2023lets]: A benchmark consisting of 500 high-difficulty competition-style mathematics problems from the MATH [hendrycksmath2021] dataset, designed to evaluate symbolic reasoning, multi-step deduction, and advanced mathematical problem-solving abilities of language models.
  • AMC23 [mathai_amc23]: A dataset derived from the 2023 American Mathematics Competition (AMC), which contains more challenging problems than GSM8K. It is commonly used to assess advanced mathematical reasoning and symbolic manipulation.

### v2

@@ 1 Introduction
Reinforcement learning (RL) has become a central paradigm for training large-scale models to exhibit complex reasoning and decision-making abilities [7, 33, 24, 2, 11, 32]. Policy optimization algorithms such as Proximal Policy Optimization (PPO) [29], Group Relative Policy Optimization (GRPO) [11, 30], and REINFORCE with leave-one-out baselines (RLOO) [1] explore different designs of reward modeling and advantage estimation to guide learning. Despite these advances, training stability remains a fundamental bottleneck, particularly in the post-training stage of large language models [6, 27, 41, 34].
Training stability is influenced not only by algorithmic design but also by choices such as learning rate schedules and baseline functions. In large-scale pre-training, adaptive or decayed learning rate schedules are routinely used to ensure stable optimization and steady convergence [3, 4, 10, 19, 9, 18]. It is therefore natural to ask whether similar benefits could be realized in reinforcement learning, where instability remains a persistent challenge [37]. Baseline design has likewise been explored as a practical tool for variance reduction, with numerous heuristics proposed across different policy optimization methods [43, 1, 35]. However, most of these strategies are empirical in nature, and the field still lacks systematic analysis that clarifies what constitutes an effective baseline and how it influences learning dynamics.
These limitations highlight the need for deeper theoretical guidance. While empirical techniques for stability are widely adopted, systematic theory for post-training policy optimization is still scarce [25, 40, 5, 39, 20]. Under mild and interpretable assumptions, rigorous analysis can illuminate the statistical properties of gradient estimators, reveal how baseline choices interact with learning dynamics, and clarify when adaptive learning rates can provably improve stability. Such theoretical insights not only fill a long-standing gap in understanding but also directly motivate algorithmic designs that bridge principled analysis with practical effectiveness.
@@ 2 Related Work
Recent studies have analyzed the training dynamics of policy optimization, especially focusing on the loss function upper bound and the resulting convergence guarantees [20, 25, 5, 39]. For example, the impact of a single-step update on the loss, along with the identification of the optimal update vector, has been analyzed [20]. Similarly, stochastic no-regret oracle frameworks have been employed to provide theoretical guarantees, leading to regret upper bounds and formal connections to online learning [5]. In parallel, classical gradient descent optimization theory has been applied to establish convergence rates under smoothness and convexity assumptions [25, 39]. Related efforts also bridge theory and practice by providing convergence results for GRPO and related algorithms, thereby supporting their empirical success [25]. Moreover, under smoothness conditions, it has been shown that the loss function admits a guaranteed decreasing rate [39]. Building on this analysis, our work establishes tighter upper bounds, provides stronger and more general convergence guarantees, and introduces principled strategies for learning rate schedules and baseline design.
@@
Policy optimization, originating from reinforcement learning, is central to shaping the reasoning capabilities of large language models and has therefore become the foundation of reinforcement learning from human feedback (RLHF) [7, 33, 24, 2, 11]. Within this framework, supervised fine-tuning (SFT) [24, 36] provides initial alignment through imitation on instruction data, PPO [29] extends this paradigm with critic-based reinforcement learning on preference-model rewards, and Direct Preference Optimization (DPO) [26] further simplifies the objective by introducing a contrastive loss that directly matches human preferences between responses. To mitigate the complexity of critic-based methods, GRPO [11, 30] replaces the learned value function with a group-based baseline defined as the average reward within each candidate set, while alternative designs such as ReMax [21], RLOO [1], and Reinforce++ [16] adopt maximum-reward, leave-one-out, or variance-reduced baselines, respectively. Building on these developments, this work presents a general formulation that unifies the above algorithms under a common framework, enabling systematic theoretical analysis and leading to the identification of a principled optimal baseline.
@@ 3.2 Related RL Algorithms
The Proximal Policy Optimization (PPO) [29] objective is formulated as
@@
where the advantage estimator $\hat{A}^{\operatorname{PPO}}$ is computed using Generalized Advantage Estimation (GAE) [28], a widely adopted variance-reduction technique that stabilizes policy-gradient training.
@@
Group Relative Policy Optimization (GRPO) [30] optimizes policies by comparing rewards among a group of sampled outputs, without relying on an explicit value or reward model. Given a query $q$, we draw $G$ outputs $\{o_{i}\}_{i=1}^{G}$ from the old policy $\pi_{\theta_{\operatorname{old}}}$ with associated rewards $r_{i}=F(q,o_{i})$. The objective is
@@
The ReMax [21] method draws inspiration from the REINFORCE with Baseline approach, where we modify the gradient estimation by incorporating a subtractive baseline value. The objective is:
@@
REINFORCE Leave-One-Out (RLOO) [1, 17] extends the REINFORCE estimator to the multi-sample setting by employing a leave-one-out baseline. Given a query $q$, we draw $G$ outputs $\{o_{i}\}_{i=1}^{G}$ from the old policy $\pi_{\theta_{\operatorname{old}}}$ with rewards $r_{i}=F(q,o_{i})$. The objective is
@@ 3.4 Assumptions
PPO [29]
GRPO [30]
ReMax [21]
RLOO [1]
@@ 4.1 Bias and Variance Analysis
Next, we analyze the covariance of the noise term $\xi(\theta)$. To simplify the expression, we define the single-sample covariance under $\pi_{\theta}(\cdot|q)$ as
$\boldsymbol{H}(\theta):=\operatorname{Var}\Big[\nabla_{\theta}\log\pi_{\theta}(o|q)\big(F(q,o)-b_{\theta}(q)\big)\Big].$ (25)
Here $o$ and $o^{\prime}$ denote two distinct samples (typically i.i.d.) from $\pi_{\theta}(\cdot|q)$. The cross-sample covariance is
$\boldsymbol{C}(\theta)=\operatorname{Cov}\Big[$ $\nabla_{\theta}\log\pi_{\theta}(o|q)\big(F(q,o)-b_{\theta}(q)\big),$
$\nabla_{\theta}\log\pi_{\theta}(o^{\prime}|q)\big(F(q,o^{\prime})-b_{\theta}(q)\big)\Big].$ (26)
Theorem 2 (Variance Expression).
@@
$\text{Var}[\xi(\theta)]=\frac{1}{N_{t}G_{t}}\boldsymbol{H}(\theta)+\frac{G_{t}-1}{N_{t}G_{t}}\boldsymbol{C}(\theta).$ (27)
@@ 4.2 Deriving an Upper Bound for the Loss Function
$\mathbb{E}[\mathscr{L}(\theta_{T})]$ $\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}$
$\quad+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}$
$\quad+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\frac{\eta_{t}^{2}}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t})).$ (29)
@@ 4.3 Optimal Learning Rate Schedule
$\eta_{t}$ $=\frac{1}{BL+B^{2}M}\cdot\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}$ (31)
$=\frac{1}{BL+B^{2}M}\cdot\frac{N_{t}\operatorname{SNR}(\theta_{t})}{1+N_{t}\operatorname{SNR}(\theta_{t})}.$ (32)
@@
$\operatorname{SNR}(\theta)$ $=\frac{\mathbb{E}\|\nabla_{\theta}[J(\theta)]\|_{2}^{2}}{\mathbb{E}[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\,\hat{A}^{\mathrm{PO}}(q,o)-\nabla_{\theta}[J(\theta)]\|_{2}^{2}]}$ (33)
$=\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta)\|_{2}^{2}}{\operatorname{tr}(\boldsymbol{H}(\theta))}.$ (34)
This theorem shows that the optimal learning rate is governed by the signal-to-noise ratio (SNR) of the gradient [40, 14].
@@
$-\sum_{t=0}^{T-1}\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{2(BL+B^{2}M)\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))\right)}.$ (35)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}=O\left(\frac{1}{\sqrt{T}}\right),$ (36)
@@ 4.4 Optimal Baseline Design
This expression reveals that the optimal baseline depends on the gradient magnitudes rather than uniform averaging, leading to the following insight, which is consistent with observations reported in prior work [12].
Takeaway 2 The baseline should not be a simple average, but a gradient-weighted one.
@@ 5 Methodology
$\hat{\eta}_{t}=\eta_{0}\frac{N_{t}\widehat{\operatorname{SNR}(\theta_{t})}}{1+N_{t}\widehat{\operatorname{SNR}(\theta_{t})}}$, (38)
@@ 6 Experiments
In this section, we present a comprehensive overview of our experimental setup, evaluation metrics, and results. We evaluate the proposed Optimal Baseline and Learning-Rate Policy Optimization (OBLR-PO) algorithm on the Qwen3-4B-Base and Qwen3-8B-Base models [38], comparing its performance against the Group Relative Policy Optimization (GRPO) baseline. All experiments are implemented with the VERL framework [31] and conducted on four H200 GPUs to enable large-scale training. The subsequent subsections detail the training configuration, evaluation protocols, and empirical results.
@@
  • OlympiadBench [13]: A benchmark of olympiad-level bilingual multimodal scientific problems, designed to evaluate high-difficulty reasoning, cross-modal understanding, and advanced problem-solving capabilities of language models.
  • GSM8K [8]: A benchmark of grade-school level math word problems, designed to measure arithmetic reasoning and step-by-step problem-solving skills of language models.
  • AIME25 [42]: A benchmark consisting of problems from the 2025 American Invitational Mathematics Examination (AIME), designed to assess advanced mathematical reasoning, precise numeric computation, and multi-step problem-solving abilities of language models.
  • MATH500 [15, 22]: A benchmark consisting of 500 high-difficulty competition-style mathematics problems from the MATH [15] dataset, designed to evaluate symbolic reasoning, multi-step deduction, and advanced mathematical problem-solving abilities of language models.
  • AMC23 [23]: A dataset derived from the 2023 American Mathematics Competition (AMC), which contains more challenging problems than GSM8K. It is commonly used to assess advanced mathematical reasoning and symbolic manipulation.